# The complexity of computing the period and the exponent of a digraph

### v1

@@ Abstract
The period of a strongly connected digraph is the greatest common divisor of the lengths of all its cycles. The period of a digraph is the least common multiple of the periods of its strongly connected components that contain at least one cycle. These notions play an important role in the theory of Markov chains and the analysis of powers of nonnegative matrices. While the time complexity of computing the period is well-understood, little is known about its space complexity. We show that the problem of computing the period of a digraph is NL-complete, even if all its cycles are contained in the same strongly connected component. However, if the digraph is strongly connected, we show that this problem becomes L-complete. For primitive digraphs (that is, strongly connected digraphs of period one), there always exists a number $m$ such that there is a path of length exactly $m$ between every two vertices. We show that computing the smallest such $m$, called the exponent of a digraph, is NL-complete. The exponent of a primitive digraph is a particular case of the index of convergence of a nonnegative matrix, which we also show to be computable in NL, and thus NL-complete.
@@ 1 Introduction
To reason about more general patterns of zero entries in powers of a nonnegative matrix, the following natural homomorphism is useful. Let $\chi(A)$ be the matrix obtained from a nonnegative matrix $A$ by replacing every strictly positive entry with 1. Denote by $\mathbb{R}_{\geq 0}$ the set of nonnegative real numbers, and by $\mathbb{B}$ the Boolean semiring $\{0,1\}$ (with usual addition and multiplication except that $1+1=1$). The mapping $\chi\colon\mathbb{R}_{\geq 0}^{n\times n}\to\mathbb{B}^{n\times n}$ is then a homomorphism from the monoid of nonnegative matrices to the monoid of matrices over $\mathbb{B}$; that is, $\chi(AB)=\chi(A)\chi(B)$ holds for all $A,B\in\mathbb{R}_{\geq 0}^{n\times n}$.
@@ A $(v_{1},v_{k})$-path (or simply a path) in a digraph $G$ is an alternating sequence
of vertices and edges of $G$. The length of this path is $k-1$. A path is a cycle if $v_{1}=v_{k}$. A digraph $G$ is called strongly connected if for every pair $u,v$ of its vertices there is a $(u,v)$-path and a $(v,u)$-path in $G$. The binary relation “for vertices $u$ and $v$, there is a $(u,v)$-path and a $(v,u)$-path” is an equivalence relation on the set of vertices, and the classes of this relation are called strongly connected components. We transfer the notions of period, index of convergence and primitivity from a matrix $A$ to its digraph $G(A)$.
It is easy to show that a digraph is primitive if and only if it is strongly connected and has period one. Moreover, if a digraph is strongly connected, then its period is the greatest common divisor of the lengths of all its cycles. If it is not strongly connected, then its period is the least common multiple of the periods of its strongly connected components that contain at least one cycle. Proofs of these statements can be found in [Ros57, Kim82]. In particular, if a digraph is primitive, then there exists $m\geq 1$ such that there is a path of length exactly $m$ between every two vertices. The index of convergence of a primitive digraph is often called its exponent of primitivity, or just exponent.
@@
Deciding if a digraph is primitive is NL-hard by a simple reduction from the problem of deciding if a digraph is strongly connected, which is NL-complete [Sip13]: it is enough to add a self-loop edge to every vertex of the input digraph. Then the period of the resulting digraph is one, and hence it is primitive if and only if the original digraph is strongly connected. This leaves open the possibility that computing the period without testing for strong connectivity might be easier. Therefore, our two main results concern the complexity of computing the period for digraphs under restrictions on their strongly connected components: we show that computing the period is NL-complete even for digraphs with a single non-trivial strongly connected component (Theorem 2.1), but that, in contrast, computing the period of strongly connected digraphs is L-complete (Theorem 3.1). The latter result involves a new relation between the period of a strongly connected digraph and reachability properties of undirected graphs (Lemma 3.3). We complement these theorems with an NL-completeness result on the exponent of primitivity (Theorem 4.1).
@@ 2 The period of a general digraph
We first show that computing the period of a digraph is NL-complete, even when it is very close to being strongly connected. A strongly connected component of a digraph is called sink if there is no path from a vertex belonging to it to a vertex not belonging to it. We call a digraph almost strongly connected if it has a unique sink strongly connected component, and every cycle belongs to this component. In this section we show the following result.
@@ 2.1 NL-algorithm
For a general (not necessarily strongly connected) digraph $G$, the period equals the least common multiple of the periods of the strongly connected components [Ros57]. Therefore, it can be exponential in $n$, and thus has to be represented in binary. Since the list of the strongly connected components of $G$ can be computed in NL, by using the algorithm from the previous paragraph, we can also compute in NL the list of their periods, say $n_{1},\ldots,n_{k}$, in their unary representation. Since NL-computable functions can be composed, to show that the overall period can be computed in NL, it suffices to show that the lcm problem can be computed in NL, where the lcm problem is to compute, in binary representation, the least common multiple of a given list of numbers represented in unary. Hence, our remaining task is to show that the lcm problem can be computed in NL. In fact, we show that it can be computed in L. The fact that the input to the lcm problem is in unary simplifies our task.
@@ 2.2 NL-hardness
We now show NL-hardness in Theorem 2.1. Given a digraph $G=(V,E)$ and two its vertices $s$ and $t$, the $(s,t)$-reachability problem asks if there exists an $(s,t)$-path in $G$. The $(s,t)$-reachability problem problem is well known to be NL-complete [Sip13]. We will make some assumptions on the digraph, which are justified by the following folklore lemma.
@@
By construction, there are now no edges incoming to $s$ and no edges outgoing from $t$. Finally, we merge vertices in such a way that that there is only one vertex $s^{\prime}\neq s$ without incoming edges and only one vertex $t^{\prime}\neq t$ without outgoing edges. ∎
We now proceed to the proof of NL-hardness in Theorem 2.1. We construct a digraph $G^{\prime}=(V^{\prime},E^{\prime})$ by modifying $G$ as follows. First, subdivide every edge of $G$ into a directed path of length two by adding a new vertex for every edge. Denote by $V^{\prime}$ the set of vertices in the obtained digraph, and by $E^{\prime}$ the set of thus obtained edges together with the edges $(s,t),(t^{\prime},t),(t,s)$. See Figure 1 for an example.
@@ 3.2 L-hardness
Figure 2: Digraphs obtained from the directed paths $v\to s\to t\to v^{\prime}$ (left) and $v\to t\to s\to v^{\prime}$ (right). Dotted vertices represent the result of subdiving edges, dashed vertices are the additional new vertex $s^{\prime}$ and the end of the path $v^{\prime}$, and dashed edges are the “special” edges going from $s^{\prime}$ and $v^{\prime}$. Depending on the order of $s$ and $t$, these “special” edges are included in different cycles and thus change the parity of the period.
@@ 4.1 NL-algorithm
Let $G=(V,E)$ be a digraph. Denote by $P$ its period, that is, the least common multiple of the periods of the strongly connected components. We want to show the first statement of Theorem 4.1, i.e., that the index of convergence of $G$ can be computed in NL. As proved in [Sch70] (see also [JyQ87, HL66]), the index of convergence is at most $|V|^{2}$. Let $0\leq K<|V|^{2}$. It suffices to show that one can determine in NL if $K$ is less than the index of convergence. This condition holds true if and only if there are vertices $s,t$ such that either
@@ 4.2 NL-hardness
To prove NL-hardness, we reduce the $(s,t)$-reachability problem in digraphs with assumptions as in Lemma 2.2 (including those on the special vertices $s^{\prime},t^{\prime}$). Let $G=(V,E)$ be such a digraph. Construct a digraph $G^{\prime}=(V^{\prime},E^{\prime})$ by adding vertices and edges to $G$ as follows. Take $V^{\prime}=V\cup U$, where $U=\{u_{1},\ldots,u_{|V|}\}$ is a set of $|V|$ fresh vertices. Thus, $|V^{\prime}|=2|V|$. Take $E^{\prime}$ to be the set $E$ together with an edge from every vertex in $V^{\prime}$ to $s$, an edge from $t$ to every vertex in $V^{\prime}$, and the edges $(t^{\prime},u_{1}),(u_{i},u_{i+1})$ for all $11\leq i\leq|V|-$ 1, and $(u_{|V|},t)$, which form a $(t^{\prime},t)$-path via $U$. This construction is illustrated in Figure 3.

### v2

@@ Abstract
The period of a strongly connected digraph is the greatest common divisor of the lengths of all its cycles. The period of a digraph is the least common multiple of the periods of its strongly connected components. These notions play an important role in the theory of Markov chains and the analysis of powers of nonnegative matrices. While the time complexity of computing the period is well-understood, little is known about its space complexity. We show that the problem of computing the period of a digraph is NL-complete, even if all its cycles are contained in the same strongly connected component. However, if the digraph is strongly connected, we show that this problem becomes L-complete. For primitive digraphs (that is, strongly connected digraphs of period one), there always exists a number $m$ such that there is a path of length exactly $m$ between every two vertices. We show that computing the smallest such $m$, called the exponent of a digraph, is NL-complete. The exponent of a primitive digraph is a particular case of the index of convergence of a nonnegative matrix, which we also show to be computable in NL, and thus NL-complete.
@@ 1 Introduction
To reason about more general patterns of zero entries in powers of a nonnegative matrix, the following natural homomorphism is useful. Let $\chi(A)$ be the matrix obtained from a nonnegative matrix $A$ by replacing every strictly positive entry with 1. Denote by $\mathbb{R}_{\geq 0}$ the set of nonnegative real numbers, and by $\mathbb{B}$ the Boolean semiring $\{0,1\}$ (with usual addition and multiplication except that $1+1=1$). The mapping $\chi\colon\mathbb{R}_{\geq 0}^{n\times n}\to\mathbb{B}^{n\times n}$ is then a homomorphism from the monoid of nonnegative matrices to the monoid of matrices over $\mathbb{B}$; that is, $\chi(A\cdot B)=\chi(A)\cdot\chi(B)$ holds for all $A,B\in\mathbb{R}_{\geq 0}^{n\times n}$. Intuitively speaking, $\chi$ agrees with matrix multiplication.
@@
Example 1.1.
Consider the nonnegative matrix
$A=\begin{pmatrix}1&1&0\\ 0&0&1\\ 0&1&0\end{pmatrix}\ \in\ \mathbb{R}_{\geq 0}^{3\times 3}\,.$
We have
$A^{2}=\begin{pmatrix}1&1&1\\ 0&1&0\\ 0&0&1\end{pmatrix}\,,\ A^{3}=\begin{pmatrix}1&2&1\\ 0&0&1\\ 0&1&0\end{pmatrix}\,,\ A^{4}=\begin{pmatrix}1&2&2\\ 0&1&0\\ 0&0&1\end{pmatrix}\ \in\ \mathbb{R}_{\geq 0}^{3\times 3}\,.$
On the other hand,
$\chi(A)=\begin{pmatrix}1&1&0\\ 0&0&1\\ 0&1&0\end{pmatrix}\,,\ \chi(A^{2})=\begin{pmatrix}1&1&1\\ 0&1&0\\ 0&0&1\end{pmatrix}\,,\ \chi(A^{3})=\begin{pmatrix}1&1&1\\ 0&0&1\\ 0&1&0\end{pmatrix}\,,\ \chi(A^{4})=\begin{pmatrix}1&1&1\\ 0&1&0\\ 0&0&1\end{pmatrix}\ \in\ \mathbb{B}^{3\times 3}\,,$
where $\chi(A^{k})=\chi(A)^{k}\in\mathbb{B}^{3\times 3}$ for all $k\geq 0$. For all $i\geq 2$ we have $\chi(A)^{i+2}=\chi(A)^{i}$. The period of $A$ is $2$, and the index of convergence of $A$ is also $2$.
@@ A $(v_{1},v_{k})$-path (or simply a path) in a digraph $G$ is an alternating sequence
of vertices and edges of $G$ with $k\geq 2$. The length of this path is $k-1$. A path is a cycle if $v_{1}=v_{k}$. A digraph $G$ is called strongly connected if for every pair $u,v$ of its vertices there are a $(u,v)$-path and a $(v,u)$-path in $G$. A maximal (by inclusion) set $S$ of vertices such that for every two vertices $u,v\in S$ there are a $(u,v)$-path and a $(v,u)$-path is called a strongly connected component. Note that in our definition, vertices that do not belong to any cycles are not in any strongly connected component. We transfer the notions of period, index of convergence and primitivity from a matrix $A$ to its digraph $G(A)$. Moreover, powering matrices over $\mathbb{B}$ has a natural interpretation in terms of digraphs: for any $k\geq 1$, the entry $(i,j)$ in $\chi(A)^{k}$ is equal to one if and only if there is an $(i,j)$-path of length $k$ in $G$.
Example 1.2.
Let $A\in\mathbb{R}_{\geq 0}^{3\times 3}$ be the matrix from Example 1.1. The digraph $G(A)$ can be visualized as follows.
111122223333
Its strongly connected components are $\{1\}$ and $\{2,3\}$. The top-right entry of $\chi(A^{4})=\chi(A)^{4}$ is the Boolean value $1$, which indicates that $G(A)$ has at least one path of length $4$ from $1$ to $3$. (In fact, there are two such paths in $G(A)$.)
It is easy to show that a digraph is primitive if and only if it is strongly connected and has period one, see e.g. [Min88, Theorem 3.3]. Moreover, if a digraph is strongly connected, then its period is the greatest common divisor of the lengths of all its cycles. If it is not strongly connected, then its period is the least common multiple of the periods of its strongly connected components (we assume that the least common multiple of an empty set is equal to one). Proofs of these statements can be found in [Ros57, Kim82]. In particular, if a digraph is primitive, then there exists $m\geq 1$ such that there is a path of length exactly $m$ between every two vertices. The index of convergence of a primitive digraph is often called its exponent of primitivity, or just exponent.
@@
Deciding if a digraph is primitive is NL-hard by a simple reduction from the problem of deciding if a digraph is strongly connected, which is NL-complete [Sip13, Problem 8.27]: it is enough to add a self-loop edge to every vertex of the input digraph. Then the period of the resulting digraph is one, and hence it is primitive if and only if the original digraph is strongly connected. This leaves open the possibility that computing the period without testing for strong connectivity might be easier. Therefore, our two main results concern the complexity of computing the period for digraphs under restrictions on their strongly connected components: we show that computing the period is NL-complete even for digraphs with a single non-trivial strongly connected component (Theorem 2.1), but that, in contrast, computing the period of strongly connected digraphs is L-complete (Theorem 3.1). The latter result involves a new relation between the period of a strongly connected digraph and reachability properties of undirected graphs (Lemma 3.3). We complement these theorems with an NL-completeness result on the exponent of primitivity (Theorem 4.1).
@@ 2 The period of a general digraph
We first show that computing the period of a digraph is NL-complete, even when it is very close to being strongly connected. A strongly connected component of a digraph is called a sink if there is no path from a vertex belonging to it to a vertex not belonging to it. We call a digraph almost strongly connected if it has a unique strongly connected component, and that strongly connected component is a sink. For example, the digraph in Figure 1 (right) is almost strongly connected but not strongly connected, as $s^{\prime}$ and five other vertices are not in the strongly connected component. In this section we show the following result.
@@ 2.1 NL-algorithm
For a general (not necessarily strongly connected) digraph $G$, the period equals the least common multiple of the periods of the strongly connected components [Ros57]. Therefore, it can be exponential in $n$, and thus has to be represented in binary. The list of the strongly connected components of $G$ can be computed in NL (without storing it on the work tape). Since NL-computable functions can be composed, by using the algorithm from the previous paragraph, we can also compute in NL the list of their periods, say $n_{1},\ldots,n_{k}$, in their unary representation. Using the compositionality of NL-computable functions again, to show that the overall period can be computed in NL, it suffices to show that the lcm problem can be computed in NL, where the lcm problem is to compute, in binary representation, the least common multiple of a given list of numbers represented in unary. Hence, our remaining task is to show that the lcm problem can be computed in NL. In fact, we show that it can be computed in L. The fact that the input to the lcm problem is in unary simplifies our task.
@@ 2.2 NL-hardness
We now show NL-hardness stated in Theorem 2.1. Given a digraph $G=(V,E)$ and two its vertices $s$ and $t$, the $(s,t)$-reachability problem asks if there exists an $(s,t)$-path in $G$. The $(s,t)$-reachability problem is well known to be NL-complete [Sip13, Theorem 8.25]. We will make some assumptions on the digraph, which are justified by the following folklore lemma.
@@
By construction, there are now no edges incoming to $s$ and no edges outgoing from $t$. Finally, we merge vertices in such a way that there is only one vertex $s^{\prime}\neq s$ without incoming edges and only one vertex $t^{\prime}\neq t$ without outgoing edges. ∎
We now proceed to the proof of NL-hardness stated in Theorem 2.1. We construct a digraph $G^{\prime}=(V^{\prime},E^{\prime})$ by modifying $G$ as follows. First, subdivide every edge of $G$ into a directed path of length two by adding a new vertex for every edge. Denote by $V^{\prime}$ the set of vertices in the obtained digraph, and by $E^{\prime}$ the set of thus obtained edges together with the edges $(s,t),(t^{\prime},t),(t,s)$. See Figure 1 for an example.
@@ 3.2 L-hardness
Figure 2: Digraphs obtained from the directed paths $v\to s\to t\to v^{\prime}$ (left) and $v\to t\to s\to v^{\prime}$ (right). Dotted vertices represent the result of subdividing edges, dashed vertices are the additional new vertex $s^{\prime}$ and the end of the path $v^{\prime}$. The three edges added at the end are dotted, and dashed edges are the “special” edges going from $s^{\prime}$ and $v^{\prime}$. Depending on the order of $s$ and $t$, these “special” edges are included in different cycles and thus change the parity of the period.
@@ 4.1 NL-algorithm
Let $G=(V,E)$ be a digraph. Denote by $P$ its period, that is, the least common multiple of the periods of the strongly connected components. We want to show the first statement of Theorem 4.1, i.e., that the index of convergence of $G$ can be computed in NL. As proved in [Sch70] (see also [JyQ87, HL66]), the index of convergence is at most $|V|^{2}$. Let $0\leq K<|V|^{2}$. It suffices to show that one can determine in NL if the index of convergence is greater than $K$. This condition is satisfied if and only if there are vertices $s,t$ such that either
@@ 4.2 NL-hardness
To prove NL-hardness, we reduce from the $(s,t)$-reachability problem in digraphs with assumptions as in Lemma 2.2 (including those on the special vertices $s^{\prime},t^{\prime}$). Let $G=(V,E)$ be such a digraph. Construct a digraph $G^{\prime}=(V^{\prime},E^{\prime})$ by adding vertices and edges to $G$ as follows. Take $V^{\prime}=V\cup U$, where $U=\{u_{1},\ldots,u_{|V|}\}$ is a set of $|V|$ fresh vertices. Thus, $|V^{\prime}|=2|V|$. Take $E^{\prime}$ to be the set $E$ together with an edge from every vertex in $V^{\prime}$ to $s$, an edge from $t$ to every vertex in $V^{\prime}$, and the edges $(t^{\prime},u_{1}),(u_{i},u_{i+1})$ for all $11\leq i\leq|V|-$ 1, and $(u_{|V|},t)$, which form a $(t^{\prime},t)$-path via $U$. This construction is illustrated in Figure 3.
@@
5 Conclusions
We have shown that the period of a digraph can be computed in NL, and distinguishing between periods one and two is already NL-hard for almost strongly connected digraphs. For strongly connected digraphs, the complexity drops to L-completeness. We have shown that the index of convergence is computable in NL, complemented by an NL-hardness result even for computing the exponent of primitive digraphs. Thus, the computational complexity of the mentioned problems is very clear. The literature contains linear-time algorithms for computing the period. One might ask whether the exponent or index of convergence can similarly be computed in linear time.